Putting the IMU on the WWW

Dear colleague:

We are writing on behalf of the International Mathematical Union to draw your attention to a new initiative of putting the IMU on the World Wide Web. The Konrad-Zuse-Zentrum für Informationstechnik in Berlin (ZIB) has been working together with the Executive Committee of the IMU to compose a “home page” for IMU. You can access this home page using any of the standard Internet tools, such as Mosaic, Netscape, etc. through the URL:

http://elib.zib-berlin.de /IMU Gopher access is possible through the following address:

gopher://elib.zib-berlin.de:70/11imu

The IMU server can be accessed via telnet as follows:

telnet elib.zib-berlin.de

Login: imu

no password required

We invite you to try this out and send us your comments.

We hope this WWW home page will serve several purposes. The first is to inform all members of the international mathematical community of what the IMU is doing. Secondly, they can find there descriptions of various programs from which they can benefit:

the IMU lectures,

exchanges with the developing countries, conferences being sponsored, etc.

Thirdly, the IMU server is also a collection of data that everyone can use to find the addresses of the main mathematical organizations of the world. We are planning to extend the scope of the IMU server so that it will become a true “home page of the world of mathematics”. You can already find a number of links to mathematical and mathematics related information offered around the world. Just look at the “Links to the Mathematical World” in the IMU server.

Fourthly, the IMU server will give everyone immediate access to the latest information on the next International Congress (ICM98 in Berlin), as well as the ability to preregister for this congress by the WWW server of ICM98 that can be “clicked” in the IMU server or can be directly accessed through the following URL:

http://elib.zib-berlin.de/ICM98

The ICM98 server contains a forms page for “preliminary preregistration”. This is not a formal registration yet. Everybody preregistered for ICM98 will be informed in the future automatically about the progress of the organization of the congress by email and will receive the final registration material etc. this way.

Please encourage your colleagues interested in ICM98 to preregister for the congress.

Whoever does not have the possibility to use the advanced Internet tools described above can send an email to the following address:

icm98@zib-berlin.de

and writing

PRELIMINARY PREREGISTRATION

into the SUBJECT line.

In the body of the mail the following information should be contained:

Last Name:

First and Middle Name:

E-mail:

Phone Number: Fax Number: Institution: Street: ZIP Code: City: Country:

Every member country of IMU is listed in the IMU server. We have one very specific request:

Can you examine the entry dealing with your country, and check whether the names and addresses are correct? We want to make every effort to keep this data collection up to date and correct. Of course we would be happy if you could provide us with additional information that we could add to the entry of your country. In particular, links to existing or planned servers of mathematical societies, institutes etc. are most welcome. The best way to inform us of any changes is to email us the corrections and suggestions to

“groetschel@ZIB-Berlin.DE”, with a copy to “imu@impa.br”.

We have a second request. We are aware of the fact that Internet communication is sometimes slow due to network overload. We would therefore be glad if some institution in your country would agree to mirror both the IMU and the ICM98 servers so that mathematicians in your country have a faster and more comfortable access. Technical details concerning this matter will be handled by the Konrad-Zuse-Zentrum für Informationstechnik in Berlin (ZIB). An institution willing to mirror the servers should contact Martin Groetschel (email: groetschel@zib-berlin.de).

Finally we have a third request. We would be happy if you could distribute the information contained in this letter to the mathematicians in your society and your country in whatever form you find appropriate (emailing a copy of this letter to interested colleagues, a note about this letter in your newsletter, etc.).

Sincerely yours,

Martin Groetschel (President of the ICM98 Organizing Committee and Vice-President of ZIB)

David Mumford (IMU President) Jacob Palis (IMU Secretary)

Addresses: Martin Groetschel Konrad-Zuse-Zentrum fuer Informationstechnik Heilbronner Str. 10 D-10711 Berlin, GERMANY Tel: +49/30/89604-210 -125 (secr) Fax: +49/30/89604-208 E-Mail: groetschel@zib-berlin.de

David Bryant Mumford Dept. of Mathematics Harvard University Science Center, One Oxford St. Cambridge, MA 02138, USA

Tel: +1 617 495-5487

Fax: +1 617 495 5132

E-Mail: mumford@math.harvard.edu

Jacob Palis Instituto Matematica Pura e Aplicada (IMPA) Estrada Dona Castorina 110 Jardim Botanico 22460-320 Rio de Janeiro, RJ, BRAZIL

Phone: +55 21-511 1749 or 21-294 9447

Fax: +55 21-512 4112 or 21-512 4115

E-Mail: jpalis@impa.br or imu@impa.br